OpenStudy (anonymous):

Find the limit of the function by using direct substitution. lim x --> pi/2 (2e^x sin x)

OpenStudy (anonymous):

I got 1. Is that correct?

hartnn (hartnn):

just plug in x =pi/2 \(\Large 2 e^{\pi/2} \sin (\pi/2)\) and sin (pi/2) = 1

OpenStudy (anonymous):

THANKS!

hartnn (hartnn):

welcome ^_^